OpenStudy (narissa):

how many solutions does this system of equation have? y=-2x+3 y=-2x-1

OpenStudy (narissa):

A.0 B.1 C.2 D.infinitley many

imqwerty (imqwerty):

from 1st equation you know that \(y=-2x+3\) putting this valur of \(y\) in the 2nd equation we get this-> \(y=-2x-1\) putting value of \(y\) from eq 1- \(-2x+3=-2x-1\) try to simplify this^

OpenStudy (narissa):

i have no idea how to simplify that :/

imqwerty (imqwerty):

try adding \(2x\) on both sides tell what u get

OpenStudy (narissa):

1 and -1

imqwerty (imqwerty):

no :| okay we had this- \(-2x+3=-2x-1\) we add 2x on both sides \(-2x+3+2x=-2x-1+2x\) we see that \(x\) gets eliminated from the whole equation and we get this- \(3=-1\) ^this is clearly not true so therefore we can conclude that whatever the value of \(x\) be it won't satisfy the equation okay? :)

OpenStudy (narissa):

oooooohhhhh ok :)

imqwerty (imqwerty):

okay now what can you conclude about the number of solutions from here?

OpenStudy (narissa):

0

imqwerty (imqwerty):

yes correct :)

OpenStudy (narissa):

thank you

imqwerty (imqwerty):

yw

OpenStudy (tkhunny):

This is an eyeball proble, The slopes are given up front. How many solutions does this system of equation have? y=-2x+3 y=-2x-1 Same Slope! This narrows it to zero or infinitely many. Different Intercept. This narrows it to zero. Done. No need to calculate anything.